Find the magnitude of the wave vector of a free electron that has the same energy as a photon with wavelength 550mm.
Answer #1

Wave vector it's magnitude is defined as wave number.

Energy of photon=hc/LAMDA = 1.24/550000= 2.254*10^(-6)ev

Wave no magnitude=wave number = 2π/LAMDA =2π/.550 =11.42